For the given 2-sets of data . Find the test value (TV)

A

= 17.4   = 17   = 2.5

B

= 14.7  = 6  = 4.8



1
Expert's answer
2021-05-13T01:15:41-0400

For A test value

"=\dfrac{17.4-17}{2.5}=\dfrac{0.4}{2.5}=1.6"


For B test value-


"=\dfrac{14.7-6}{4.8}=\dfrac{8.7}{4.8}=1.5125"